# Canvil: Designerly Adaptation for LLM-Powered User Experiences

### v1

@@ 6.3.1. Integrating Adaptation into Design Practice
Designers saw great value in engaging with adaptation and saw direct paths to application in their own design practice beyond the study. P10, who has extensive experience working on AI-powered products deployed to users worldwide, said that adaptation has traditionally been difficult: “We had teams of people training models and nudging the technology to align with [user personas].” Having experienced adaptation of LLMs via Canvil, they shared that “I can see this being something that would enable [adaptation] to actually happen in a way that’s much more practical than taking Python classes, which I’ve done.” For P13, who works on a product with enterprise and consumer versions, adaptation can help set more detailed product requirements where necessary: “I could see [adaptation] being really helpful, being able to prototype and build separate generative models for enterprise customers and then to work more closely with them to tune and add requirements.” P16 considered adaptation to be a useful exercise for user empathy: “for [Audience Setting] I just tried to put myself in the mind of someone who’s using this and say, I’m really good at following recipes, but have a variety of dietary preferences and restrictions.”
Some participants pointed out that designers should be prepared to carefully navigate challenges of adaptation that may arise in practice. P11 reflected on the balance between specificity through adaptation and breadth of audience: “I wanted to avoid bias in these [models]. There’s that balance between, having the profile of the model be specific enough so that it’s relatable […] and making technology for everybody. If you try to make it too generic, it’s just gonna be bland. So we have to be careful.” P15 pointed out that adaptation may not account for all user needs and preferences: “in terms of the normative sense, it would be fantastic. But feasibly though, there are so many different things that you need to take into account, like even length of language.” These complexities highlight the broader social and ethical factors surrounding designerly adaptation to confront as designers integrate it into their own practice.
6.3.2. Designerly vs. End-User Adaptation
Prior work found that designers were naturally inclined to provide end-users with more controls for customization of their AI-powered user experiences (Feng and McDonald, 2023). Some designers also did the same in our study and reflected on the interplay between designerly and end-user adaptation. P15, who discussed the complex factors of adaptation in Section 6.3.1, said they envisioned a more feasible approach to be designers delegating adaptation to end-users past a certain point: “If we could put in a Venn diagram of all of the different needs from all over the world, seeing what the more central options are, the really interesting piece here is having users be able to make that [adaptation] decision themselves.” P10, after interacting with Canvil, started thinking about similar tools for end-user adaptation: “I thought [Canvil] could be a tool, yes, for us content designers, but could we adopt this for individual users so they could tweak [the model]?” However, they acknowledged that granting end-users complete control over adaptation may not be ideal either, and designers are responsible for setting certain constraints:
“In my experience with human beings, you don’t want to give them a blue sky. You wanna give them things they can create on their own and feel like they’re in control. I could see us using [Canvil] to create something like 5 or 6 different [model] personalities that users around the world could leverage.” [P10]
P11 also found the idea of balancing designerly and end-user adaptation appealing, as they were concerned about model biases (see Section 6.3.1). They proposed having “a preference or a setting or a dial based on your regional settings where [LLM behavior] could be modulated.” With the rapid rise of LLM-powered user experiences, we invite practitioners and researchers to more deeply reason about who can be responsible for adaptation and when adaptation should happen.
6.3.3. Areas of Additional Support
@@
Designers’ also identified several areas for improvement to better support designerly adaptation. For one, designers lacked a clear mental model of how the fields of Canvil would impact the adaptation outcome, including how much detail is required in authoring the prompt. As a result, some designers wrote long, detailed instructions, while others kept their Canvils sparse, and some were pleasantly surprised by how well the model handled minimal instructions. P6 further noted that the stochastic nature of the model was particularly challenging to work with when writing outputs to their designs: “every time I generated the [response], I felt that they were different every time, so it was not easy to predict what the next [response] would look like.” We note that current LLMs are known to be challenging to be precisely controlled through prompting, and there is a lack of transparency into (or even an established understanding of) how prompting impacts LLM behaviors. However, these are actively researched areas which can help improve the mental model of Canvil’s users and its general effectiveness.
@@
Additionally, some designers wanted to engage in finer-grained experimentation and iteration by only focusing on a specific field at once. P7, who wanted to iterate more on the Model Profile field, wondered if there was a way to “decrease the size [of the other fields] and expand to view everything.” We envision a modular future version of the tool where each field can be separated, such that a user can mix-and-match different fields to form a complete system prompt. However, before that, we may need to address the precise mapping between each field (and potential overlaps between them, as noted by some participants) to adaptation outcomes as discussed above. As mentioned in Section 4.3, our fields present just one possible structure for system prompts. The fields can perhaps be reconfigured to reduce potential overlap, or even dynamically generated based on the design task.
@@ 7.1. A Workflow for Designerly Adaptation
Drawing on the results from both our formative and design study, we now propose a workflow for designerly adaptation. Our proposed workflow consists of four steps:
@@
    Understand deployment context through user research. To orient adaptation, it is imperative to first understand the context in which users will interact with the LLM-powered system. This includes users’ goals, needs, and pain points, along with customs and values that may affect their use of the technology. Those with expertise in user research methodologies should lead the execution of this step.
@@
    Embed user research insights into system prompts, learning from examples where possible. This carves out a direct path for user research to impact model behavior, as we observed in Section 6.1.1. Thanks to the collaborative nature of many modern design tools (Feng et al., 2023c), there may be example prompts written by other designers available for reference, or templates to use as a starting point (Section 6.2.1). By leveraging collaborative affordances to share knowledge, designers can enhance their efficiency at adaptation over time.
@@
    Co-evolve interface design and prompts. As observed in Section 6.1.2, adaptation can supply new inspiration for UI designs and affordances. On the other hand, designers also tinkered with prompts to coax the model into providing outputs that fit into the constraints laid out by the designs being explored. We see the co-evolution of designs and prompts as a promising path forward, in which iterative tinkering with system prompts for adaptation shapes design choices, and vice versa.
@@
    Share designs and adaptation efforts with the broader team. Showcasing in-progress work through design critiques is already a part of the design process (Feng et al., 2023b). In our study, we found that sharing Canvils helped envision new collaborative workflows with other designers, as well as communicating their perspectives and negotiating with technical stakeholders (Section 6.2.1). Following from Step 3, we thus believe that sharing designs alongside adaptation is critical for involving designers and leveraging their expertise for responsible development and deployment of LLM-powered systems.
@@
7.2. Towards Collaborative Tinkering With AI
Our work highlights the promises of not only empowering designers to tinker with LLMs as a design material, but also collaboratively doing so. During our design study, designers proposed various collaborative workflows for system prompting with Canvil, including a divide-and-conquer approach to gather expertise distributed across the team and a tinker-and-unite approach where each designer tinkers independently before merging the results into a team version (Section 6.2.1). Additionally, designers benefited from seeing and learning from others’ Canvils and commented on more organized knowledge sharing and version management afforded by using Canvil in Figma’s multiplayer canvas environment (Section 6.2.2).
Despite promising opportunities, many tools that lower the barrier for AI tinkering (e.g., (Lobe.ai, 2021; Liner.ai, 2022; Carney et al., 2020; Subramonyam et al., 2021a)) were designed with individualized workflows in mind. Some, like Google’s Teachable Machine (Carney et al., 2020), offer social features such as project galleries to showcase example projects, but affordances for collaboration remain limited. One reason may be technical constraints surrounding the collaborative updating of computationally heavy stages of large-scale training pipelines. The shift towards using large pre-trained models, and the ability to interact with these models via natural language, helps accelerate model adaptation and tightens feedback loops during tinkering. Consequently, novel interfaces, such as node-based editors, (e.g., (Arawjo et al., 2023; Suh et al., 2023; Wu et al., 2022b; Angert et al., 2023)), enable users to decompose large tasks by “chaining” together smaller prompt submodules. Node-based editors are conceptually appealing for collaboration (Wu et al., 2022b)—collaborators may distribute work according to expertise and nature of the task within the submodule, and easily chain together their work. Few node-based editors, however, support multiplayer collaboration. We see significant potential in extending these tools along a collaborative dimension.
Besides technical constraints discussed, it can simply be the case that robust collaborative experiences are generally challenging to implement. We circumvented this in our work by building on top of the Figma API, which allows us to leverage Figma’s built-in collaborative features by default. We thus encourage researchers and practitioners who wish to build collaborative tools to take advantage of existing collaborative platforms’ developer APIs where possible, especially as these APIs become more richly featured.
@@ 7.3. Materiality and Sociomateriality of LLMs
Throughout this work, we depict LLMs as a design material to use materiality as a means of understanding, shaping, and applying this technology to tackle design problems. Yet, in collaborative settings, materiality’s implications extend beyond the individual. Scholars use the term sociomateriality in recognition of materiality’s tendency to shape, and be shaped by, organizational practices typically constituted as “social” (e.g., decision-making, strategy formulation) (Orlikowski, 2007; Leonardi, 2012). For example, Orlikowski observed that the issuance of BlackBerry devices within a company led employees to obsessively check for new messages and send immediate responses (Orlikowski, 2007). The BlackBerrys’ material properties—in this case, being able to receive and send messages on-the-go—reconfigured employees’ social practices, which in turn shifted how they think and act with the technology.
@@
In our design study, designerly adaptation allowed for exploration of LLMs’ material properties through tinkering. As a result, designers proposed various UI affordances for enhancing users’ interaction with the LLM-powered meal planner (Section 6.1.2) and adjusted the model’s guardrails depending on differing dietary restrictions between user groups. However, grappling with the materiality of LLMs also demonstrated potential to reconfigure social practices within product teams. For example, designers informed us that they saw new avenues of collaboration upon interaction with Canvil, which included “handing off” system prompts to data scientists by exporting prompts out of Figma, working with PMs to inform prompts with product requirements, and having a more concrete artifact to communicate desired tweaks to engineers (Section 6.2.3). Moreover, designerly adaptation itself may only be one piece of the broader adaptation puzzle. Certain capabilities such as storing external knowledge in model weights through fine-tuning cannot easily be unlocked through—and may even a prerequisite for—designerly adaptation, paving the way for processes and artifacts that fundamentally differ from those used in tradition design handoffs (Leiva et al., 2019; Feng and Zhang, 2022). This potential reconfiguration of teams’ collaborative dynamics establishes designerly adaptation as not only a means of exploring LLMs’ materiality, but also a sociomaterial practice.
@@
The implications of this observation are twofold. First, sociomateriality argues that reconfigurations of collaborative practices upon interaction with a prominent new technology are inevitable (Orlikowski, 2007; Leonardi, 2012). At the time of writing, frenzied excitement over LLM capabilities has launched an industry-wide race to embed them into products and product suites (Parnin et al., 2023). There is much yet to be discovered about shifts in organizational practices and the emergence of new ones in the midst of this race, so researchers in CSCW and organizational science should be attuned to emergent challenges. Second, addressing these challenges for practitioners and providing probes for researchers to study them may require new processes and tools. Canvil is an early example of such a tool, but more are needed to tackle the multiplicity of open questions in today’s rapidly evolving AI landscape.
@@ 8. Limitations and Future Work
Our design study, conducted in Figma, aimed to mirror real-world design activities (Ball, 2005), yet concerns about ecological validity may persist. User research in practice may differ in presentation and detail than in our study setup, which can impact how designers perform research-informed adaptation. Feasto, the fictitious app in our design study, applied LLMs to the universally relatable topic of food. Designers in domains with fewer broadly-shared experiences (e.g., accessibility), may require deeper collaboration with domain experts and thus face workflow complexities not accounted for in our study. A potential direction for future work, then, is longitudinal studies that use contextual inquiry to observe product teams throughout a full development lifecycle of an LLM-powered feature, focusing on key adaptation decision-making processes and the practitioners involved.
Our study, like any study that uses a probe, has results contingent on our probe’s features. For example, designers’ interaction with the Main Form—designed with recommended practices for system prompting (Microsoft, 2023; OpenAI, 2023c)—shaped their approach to adaptation, and changes to the form could impact existing results and reveal new ones. We mitigate this by presenting findings not tied to the Main Form, nor our design study task. Future research can build probes for designerly adaptation, distinct from Canvil, to explore agreements with and divergences from our results.
Designers in our study also provided feedback on Canvil that we can integrate into future work. These include breaking down the Main Form into sub-Canvils and linking them together to assemble system prompts, offering support for multimodal models (e.g., GPT-4V [^12] At the time of our study, GPT-4V was not yet widely accessible via APIs, so the most advanced model available for use in Canvil was GPT-4.), and text formatting controls for model outputs. Finally, as mentioned in Section 4.4, Canvil may also be used in FigJam. Future research could explore adaptation in early design stages, like brainstorming and ideation, through studies in FigJam.

### v2

@@ 6.2.1. Designers Were Receptive to Designerly Adaptation but also Recognized its Limits
Designers saw great value in engaging with adaptation and saw direct paths to application in their own design practice beyond the study. P10, who has extensive experience working on AI-powered products deployed to users worldwide, said that working with AI as a design material has traditionally been difficult: “We had teams of people training models and nudging the technology to align with [user personas].” Having experienced LLM adaptation via Canvil, they shared that “I can see this being something that would enable [adaptation] to actually happen in a way that’s much more practical than taking Python classes, which I’ve done.” For P13, who works on a product with enterprise and consumer versions, adaptation can help set more detailed product requirements where necessary: “I could see [adaptation] being really helpful, being able to prototype and build separate generative models for enterprise customers and then to work more closely with them to tune and add requirements.” P16 considered adaptation to be a useful exercise for user empathy: “for [Audience Setting] I just tried to put myself in the mind of someone who’s using this and say, I’m really good at following recipes, but have a variety of dietary preferences and restrictions.”
Designers also recognized that designerly adaptation is only one piece of the larger puzzle when it comes to defining an LLM’s behavior in a production-ready product. Some changes desired by designers, such as factual grounding and connecting to external knowledge bases, may still need engineered by technical teams. Both P12 and P14, for example, wanted Canvil to be able to connect to custom models built and maintained by their own teams, as the behavior of default OpenAI models may be too generic to use as a suitable design material. P14 comments: “I think linking to the vanilla ChatGPT model isn’t specific enough. If there is a way to somehow have Canvil link to the devs’ [models] while I’m designing and I’m able to see what type of experience our users are getting, I think that would be super helpful.” While we currently offer only a limited number OpenAI models to use with Canvil, custom models may be added with lightweight engineering as the widget is communicating with the model via API endpoints. This would also Canvil to integrate technical advancements into design processes.
6.2.2. Canvil Afforded Collaboration and Knowledge Sharing Within the Design Process
Many designers commented on how Canvil afforded collaboration (DG4), a much-welcomed benefit that extends beyond the context of our study. P5 appreciated that they could leverage Figma’s built-in collaboration features: “If I see my fellow designer’s Canvil, and if I want to change something then and there, I can drop a comment on it just like a normal component in Figma, which we’ve been doing in our everyday design work.” P10 shared that their team relies on a spreadsheet to keep track of model adaptation efforts and working in a canvas environment would be a significant improvement.
In particular, designers saw potential for Canvil to facilitate more effective knowledge sharing about LLM adaptation across design and non-design teams. P10 commented that: “You could have a master Canvil and then you could make copies, and [others] can then do their own interpretations.” P11 agreed and added that seeing iterations on a canvas can help find inspiration in others’ work: “I think anytime you line up different iterations together, you notice: ohh that person, did you know they had that approach? That’s a good idea. And I’m gonna try that over here. I think it really is an aid to experimentation.” We observed other instances of knowledge sharing as well when designers viewed others’ Canvils. For example, P4 noted that they were inclined to consider more dimensions for their Main Form after seeing P6’s work: “[looking at P6’s Canvil] made me thinking about timers and Hindi slang.” Some also realized new capabilities of LLMs by looking at others’ Main Form and outputs—P1 shared that “I didn’t realize at first that you can actually make the [LLM] generate multiple recipes, just like what P2 did.” P1’s groupmates (P2 and P3) were also intrigued when they saw that P1 had used one Canvil to generate example inputs for another Canvil to use. These social and collaborative affordances differentiate Canvil from prior systems for empowering designers to tinker with AI (Subramonyam et al., 2021a; Petridis et al., 2023b; Carney et al., 2020).
Designers were also excited by the use of Canvil as a boundary object (Star and Griesemer, 1989) to collaborate on adaptation across domain boundaries. P2 shared that they “can see from a product manager’s perspective that they would love to play around with [Canvil], and it would probably help the designer find more common ground because they both used a similar tool.” For deepening collaboration with data scientists and other stakeholders, P11 commented that the “copy to clipboard” feature in Generate Panel (Section 4.2.3) can allow for direct handoff of model behavioral specifications: “[The button] is a way of exporting that so a data scientist can come along and say sure, let me plug that into code.” Even if designers were unable to achieve the desired model behavior themselves, P17 thought Canvil was helpful in specifying desired changes: “If I want the model to respond in this way versus that way, just having something tangible to show engineering partners where the tweak would need to be, would be helpful. And I think [that’s] obviously easier with Canvil.”
6.2.3. Desired Areas of Additional Support
@@
Designers identified several areas of improvement to better support designerly adaptation. For one, designers lacked a clear mental model of how the fields of Canvil would impact the adaptation outcome, including how much detail is required in authoring the specifications. As a result, some designers wrote long, detailed instructions, while others kept their Canvils sparse, and some were pleasantly surprised by how well the model handled minimal instructions. P6 further noted that the stochastic nature of the model was particularly challenging to work with when writing outputs to their designs: “every time I generated the [response], I felt that they were different every time, so it was not easy to predict what the next [response] would look like.” We note that current LLMs are known to be challenging to be precisely controlled through natural language, and there is a lack of transparency into (or even an established understanding of) how natural language instructions influence LLM behaviors. However, this is an area of active research which can help improve Canvil users’ mental models and the tool’s general utility. For example, new developments in using sparse autoencoders to surface human-interpretable features in LLMs as novel controls (Templeton et al., 2024; Gao et al., 2024) may be integrated into Canvil to help designers steer models using specific features relevant to their design problem.
@@
Additionally, some designers wanted to engage in finer-grained experimentation and iteration by only focusing on a specific field in the Main Form. P7, who wanted to iterate more on the Model Profile field, wondered if there was a way to “decrease the size [of the other fields] and expand to view everything.” We envision a modular future version of Canvil where each field can be separated, such that a user can mix-and-match different fields. However, before that, we may need to address the precise mapping between each field (and potential overlaps between them, as noted by some participants) to adaptation outcomes as discussed above. As mentioned in Section 4.2, our fields present just one possible structure for model behavioral specifications via design requirements. The fields can perhaps be reconfigured to reduce potential overlap, or even dynamically generated based on the design task.
@@ 7.1. A Workflow for Designerly Adaptation
Drawing on the results from both our formative and design studies, we now propose a workflow for designerly adaptation to concretely illustrate how it might be used as a design code. We intentionally make this workflow agnostic to the specific LLM adaptation technique—if a new technique provides finer control over model behavior and is accessible to generalist audiences (e.g., feature clamping (Templeton et al., 2024)), the workflow can incorporate that technique. Our proposed workflow consists of four steps:
@@
    Understand deployment context through user research. To orient adaptation and establish user requirements, it is imperative to first understand the context in which users will interact with the LLM-powered application. This includes users’ goals, needs, and pain points, along with customs and values that may affect their use of the technology. This step should ideally be led by those with expertise in user research methodologies.
@@
    Translate user requirements into model behavior, learning from examples where possible. Appropriate tooling can carve out a direct path for user research to define design requirements that impact model behavior, as we observed in Section 6.1.2. Thanks to the collaborative nature of many modern design tools (Feng et al., 2023c), there may be example adaptation attempts by other designers available for reference, or templates to use as a starting point (Section 6.2.2). By leveraging collaborative affordances to share knowledge, designers can more quickly familiarize themselves with the new design material.
@@
    Co-evolve designs and model behavior. As observed in Section 6.1.1, model adaptation can supply new inspiration for UX designs and UI affordances. On the other hand, designers also tinkered with the Main Form to steer the model towards providing outputs that fit into the constraints laid out by existing designs. We see the co-evolution of designs and prompts as a promising path forward, in which iterative tinkering with adaptation approaches shapes design decisions, and vice versa.
@@
    Share adaptation efforts with the broader team. Showcasing in-progress work through design critiques is already a part of the design process (Feng et al., 2023b). In our study, we found that sharing Canvils helped envision new collaborative workflows with other designers, as well as communicating their perspectives and negotiating with technical stakeholders (Section 6.2.2). We thus believe that creating sharable artifacts that depict adaptation efforts—i.e., translational knowledge from the use of a design code—is an integral part of designerly adaptation.
@@
7.2. Towards Design Codes for Collaborative Material Exploration of AI
Our work highlights the promises of not only empowering designers to explore LLMs as a novel design material, but also collaboratively doing so. For example, designers thought they benefited from seeing and learning from others’ Canvils and commented on more organized knowledge sharing and version management afforded by using Canvil in Figma’s multiplayer canvas (Section 6.2.2). While many tools from prior work lower the barrier for exploring AI’s—not just LLMs’—material properties (e.g., (Lobe.ai, 2021; Liner.ai, 2022; Carney et al., 2020; Subramonyam et al., 2021a)), few offer collaborative affordances. To extend existing tools along a collaborative dimension, we encourage a shift from solely focusing on new tools to design codes which a tool can help operationalize.
Because design codes entangle underlying work processes with material exploration (Cross, 1982), they urge tool builders to consider how tools enabling this exploration can integrate with existing collaborative workflows. Thus, in our work, we developed Canvil as a Figma widget to seamlessly integrate into Figma canvases where designers already have well-established collaborative practices (Feng et al., 2023c). Developing a tool designing for collaborative workflows early on, and only later highlighting its collaborative potential, may not be sufficient to make it truly collaborative. For example, node-based editors for steering LLMs (e.g., (Arawjo et al., 2024; Suh et al., 2023; Wu et al., 2022b; Angert et al., 2023)) are conceptually appealing for multiplayer collaboration (Wu et al., 2022b), but few support it in practice. When embedding collaboration into these tools, one may consider how the tool can entangle existing collaborative processes—which may vary significantly from community to community—with LLMs. For example, understanding how development teams share prompts and model outputs may provide valuable insights for a more collaborative ChainForge (Arawjo et al., 2024), while examining social practices of the p5 creative coding community can inform collaborative features for Spellburst (Angert et al., 2023).
We note that robust collaborative experiences, particularly real-time ones, can be challenging to implement. We tackled this in our work by building on top of the Figma API, which allows us to leverage Figma’s built-in collaborative features by default. We thus encourage researchers and practitioners who wish to build collaborative tools to take advantage of existing collaborative platforms’ developer APIs where possible, especially as these APIs become more richly featured.
@@ 7.3. Materiality and Sociomateriality of LLMs
While Cross, in his original 1982 essay (Cross, 1982), discussed design codes in the context of design education, these codes have notable sociomaterial properties when used in collaborative and organizational settings. Designerly adaptation is no exception. Scholars use the term sociomateriality in recognition of materiality’s tendency to shape, and be shaped by, organizational practices typically constituted as “social” (e.g., decision-making, strategy formulation) (Orlikowski, 2007; Leonardi, 2012). For example, Orlikowski observed that the issuance of BlackBerry devices within a company led employees to obsessively check for new messages and send immediate responses (Orlikowski, 2007). The BlackBerry’s material properties—in this case, being able to receive and send messages on-the-go—reconfigured employees’ social practices, which in turn shifted how they think and act with the technology.
@@
In our design study, designerly adaptation allowed for interactive exploration of LLMs’ material properties. Grappling with these properties not only allowed designers to reason about UX improvements (Section 6.1.1), but also demonstrated potential to reconfigure social practices within product teams. For example, designers informed us that they saw new avenues of collaboration with non-design stakeholders, which included working with PMs to inform prompts with design requirements, and having a more concrete artifact to communicate desired model changes to engineers and data scientists (Section 6.2.2). Moreover, designerly adaptation may only be one piece of the broader puzzle when it comes to transforming LLMs’ materiality at a more fundamental level. Certain capabilities such as storing external knowledge in model weights through fine-tuning cannot easily be unlocked through—and may even a prerequisite for—efforts of teams working at the model level. This potential reconfiguration of desigers’ collaborative when working with LLMs establishes designerly adaptation as not only a design code, but also a sociomaterial practice.
@@
The implications of this observation are twofold. First, sociomateriality argues that reconfigurations of collaborative practices upon interaction with a prominent new technology are inevitable (Orlikowski, 2007; Leonardi, 2012). At the time of writing, frenzied excitement over LLM capabilities has launched an industry-wide race to embed them into products and product suites (Parnin et al., 2023). There is much yet to be discovered about shifts in organizational practices and the emergence of new ones in the midst of this race, so researchers in CSCW and organizational science should be attuned to emergent challenges. Second, studying and addressing these challenges may require new processes and tools. Canvil is an early example of such a tool, but more are needed to tackle the multiplicity of open questions in a rapidly evolving AI landscape.
@@ 8. Limitations and Future Work
Our design study, conducted in Figma, aimed to mirror real-world design activities (Ball, 2005), but concerns about ecological validity may remain. User research in practice may differ in presentation and detail than in our study setup, which can impact how designers perform research-informed adaptation. Feasto, the fictitious app in our design study, applied LLMs to the universally relatable topic of food. Designers in domains with fewer broadly-shared experiences (e.g., accessibility), may require deeper collaboration with domain experts and thus face workflow complexities not accounted for in our study. A potential direction for future work, then, is longitudinal studies that use contextual inquiry to observe product teams throughout a full development lifecycle of an LLM-powered feature, focusing on key adaptation decision-making processes and the practitioners involved.
Our study, like any study that uses a probe, has results contingent on our probe’s features. For example, designers’ interaction with the Main Form—designed with recommended practices for defining model behavior using natural language (Microsoft, 2023; OpenAI, 2023c; Anthropic, 2024)—shaped their approach to adaptation, and changes to the form could impact existing results and reveal new ones. We mitigate this by presenting findings not tied to the Main Form, nor our design study task. Future research can build probes for designerly adaptation, distinct from Canvil, to explore agreements with and divergences from our results.
Designers in our study also provided feedback on Canvil that we can integrate into future work. These include breaking down the Main Form into sub-Canvils and linking them together to assemble model behavioral specifications, image generation with multimodal models, and text formatting controls for model outputs. Finally, as mentioned in Section 4.3, Canvil may also be used in FigJam. Future research could explore adaptation in early design stages, like brainstorming and ideation, through studies in FigJam.